# Following the Autoregressive Nature of LLM Embeddings via Compression and Alignment

### v2

@@ 4.2 Main Results
Models without contrastive training are divided into two categories. The first is our own fair baseline model, which performs significantly worse than AutoRegEmbed, with an average performance 20% lower. This highlights the difficulty of untrained LLMs in directly generating high-quality embeddings. While some methods enhance the base model’s embeddings through prompt optimization, their improvements remain limited—even on a 13B-parameter model—and come with significant additional reasoning costs. For instance, Echo requires processing text twice to mitigate unidirectional attention limitations, MetaEOL aggregates embeddings from eight different instructions, and GenEOL relies on ChatGPT to generate up to 32 text variants before aggregating their embeddings. These additional reasoning steps severely constrain their practical implementation.
@@
LLM2Vec enhances existing LLMs using an unsupervised contrastive learning approach similar to SimCSE, leading to significant performance gains. Compared to the base model, the unsupervised version of LLM2Vec improves performance by over 15%. Although it utilizes almost 160,000 data samples, its performance remains 4.5% lower than AutoRegEmbed, demonstrating its lower efficiency.
@@
Supervised contrastive learning is the mainstream approach for building high-quality embedding models. We first compared top-performing methods that once ranked on the MTEB leaderbord (with the time they reached SOTA in brackets), including NV-Embed (2024.08), SFR-Embedding-2_R (2024.02), gte-Qwen2-7B-instruct (2024.06), and LLM2Vec (2024.05). In terms of model performance, AutoRegEmbed outperforms most SOTA methods on ten STS datasets, trailing only 0.2% behind the best version of LLM2Vec. From a data efficiency perspective, AutoRegEmbed achieves performance comparable to the previous SOTA models with just 66,382 training samples, whereas the latter requires tens of millions of triplets to reach peak performance. Additionally, previous SOTA models employ multi-task learning (e.g., retrieval and clustering), whose impact on STS performance remains unclear. To ensure a fair comparison, we use single-task contrastive learning as a baseline. Unlike traditional contrastive learning, AutoRegEmbed does not rely on in-batch negative samples. So we add two baselines to single-task contrastive learning that also exclude the in-batch negative sample strategy. As shown in Table 1, even under identical training data, AutoRegEmbed outperforms four different single-task contrastive learning, further validating its effectiveness.
@@ 4.3 Ablation Study
Figure 3: We evaluate the learning efficiency of our method against traditional contrastive learning on 10 STS datasets, comparing their performance under the same number of samples (left figure) and same number of epochs (right figure). Further details are provided in Appendix A.
To verify the effectiveness of AutoRegEmbed, we conducted an ablation study. First, we removed Conditional Distribution Alignment to evaluate its impact on model performance. Second, since Equation LABEL:eq7 was derived empirically in our previous work, we tested different variants of this equation to confirm that it remains the optimal choice. Different variants of Equation LABEL:eq7 include Log_sigmoid, which maps similarity to a logarithmic scale for integration with the exponential function e, as well as KL divergence and JS divergence, which quantify the distance between the conditional probability distributions of positive and negative sample embeddings in distinct ways. The specific equations are provided in Appendix B.
@@
Table 2 presents the ablation results. The experiments on different tasks indicate that Conditional Distribution Alignment improves performance by 9.17%, while Information Compression contributes a 16.99% improvement, demonstrating the effectiveness of both tasks. Additionally, experiments on variants of Equation LABEL:eq7 reveal that, although using a logarithmic scale for similarity and employing KL or JS divergence to measure distribution distance are more intuitive approaches, they do not surpass the performance of the original loss function in Equation LABEL:eq7. Thus, Equation LABEL:eq7 can be regarded as a more effective loss function.
@@ 4.4 Learning Efficiency
To verify that AutoRegEmbed is better suited for LLMs, we compare its performance with four contrastive learning baselines under the same training data and the same number of epochs, as shown in Figure 3. The left figure in Figure 3 shows that as the training data increases, the performance of both AutoRegEmbed and other contrastive learning methods improves, but AutoRegEmbed exhibits the fastest growth. Notably, with just 15,000 samples, AutoRegEmbed already surpasses the maximum performance of other contrastive learning models. The right figure demonstrates that as the number of epochs increases, AutoRegEmbed also improves at the fastest rate. These results indicate that AutoRegEmbed significantly outperforms the baseline models in learning efficiency.
@@ 6 Limitations
To mitigate this risk, we recommend that users of AutoRegEmbed ensure that the training data is carefully curated, and ideally, filtered for harmful content. By using safe and ethically sourced data, the model’s potential for propagating bias or harm can be minimized. Additionally, users should be cautious when applying AutoRegEmbed to sensitive domains, where the generation of unsafe or biased embeddings could have significant consequences.
@@ Appendix A Implementation details
For the information compression task, we set the learning rate to 2e-5, the batch size to 32, and train for 2 epoch. To represent the semantics of the input, we use 5 compressed tokens. For the conditional distribution alignment task, the learning rate is set to 5e-6, with a batch size of 32 and 4 epochs. The temperature parameters t and b are both set to 0.1. For the above two tasks, we set the maximum token length of context, instruction, and target to 512. Furthermore, we employ the bfloat16 format, enable FlashAttention 2, and train on four A100-80G GPUs with DeepSpeed and Zero-2. The information compression task takes 20 minutes, while the conditional distribution alignment task, involving 50,000 samples, takes approximately 1 hour.
@@ Appendix B Variants of Equation LABEL:eq7
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{-}|e_{q,I_{\mathrm{next}}})}).$
@@
$S_{2}(q,d^{-})=-\sigma(\frac{1}{T}\sum_{t=1}^{T}\mathrm{KL}(p_{\theta_{E}}(d^{+}_{t}|d^{-}_{<t},e_{d^{-},I_{\mathrm{self}}}),$
@@
$S_{2}(q,d^{-})=-\sigma(\frac{1}{T}\sum_{t=1}^{T}\mathrm{JS}(p_{\theta_{E}}(d^{+}_{t}|d^{-}_{<t},e_{d^{-},I_{\mathrm{self}}}),$

### v3

@@ 4.2 Main Results
Models without contrastive training are divided into two categories. The first is our own fair baseline model, which performs significantly worse than AutoRegEmbed, with an average performance 20% lower. This highlights the difficulty of untrained LLMs in directly generating high-quality embeddings. While some methods enhance the base model’s embeddings through prompt optimization, their improvements remain limited—even on a 13B-parameter model—and come with significant additional reasoning costs.
@@
LLM2Vec enhances existing LLMs using an unsupervised contrastive learning approach similar to SimCSE, leading to significant performance gains. Compared to the base model, the unsupervised version of LLM2Vec improves performance by over 15%. Although it utilizes almost 160,000 data samples, its performance remains 4.74% lower than AutoRegEmbed, demonstrating its lower efficiency.
@@
Supervised contrastive learning is the mainstream approach for building high-quality embedding models. We first compared SOTA methods that once ranked on the MTEB leaderbord, including NV-Embed, SFR-Embedding-2_R, gte-Qwen2-7B-instruct, and LLM2Vec. In terms of performance, AutoRegEmbed surpasses all leading methods across 10 STS datasets, outperforming LLM2Vec by a margin of 0.58. From a data efficiency perspective, AutoRegEmbed achieves performance comparable to the previous SOTA models with just 66,382 training samples, whereas the latter requires tens of millions of triplets to reach peak performance. Additionally, previous SOTA models employ multi-task learning (e.g., retrieval and clustering), whose impact on STS performance remains unclear. To ensure a fair comparison, we use single-task contrastive learning as a baseline. Unlike traditional contrastive learning, AutoRegEmbed does not rely on in-batch negative samples. So we add two baselines to single-task contrastive learning that also exclude the in-batch negative sample strategy. As shown in Table 1, even under identical training data, AutoRegEmbed outperforms four different single-task contrastive learning, further validating its effectiveness.
4.3 Performance on retrieval tasks
Table 2 summarizes the performance of various baseline methods and AutoRegEmbed across three retrieval datasets. On MS MARCO, AutoRegEmbed outperforms most prior state-of-the-art (SOTA) models and ranks second only to gte-Qwen2-7B-instruct. It also achieves competitive results on the other two datasets, NFcorpus and SCIDOCS. This outcome is expected, as AutoRegEmbed is trained solely on MS MARCO, whereas previous SOTA models are trained on large-scale datasets spanning diverse distributions, which contributes to their strong generalization on NFcorpus and SCIDOCS.
Notably, AutoRegEmbed consistently outperforms LLaMA2-inbatch-M, which is trained on the same data, across all three datasets. This further confirms the effectiveness of our method in retrieval tasks, even without large-scale multi-domain training.
Method MS MARCO NFcorpus SCIDOCS
LLM2Vec(Unsupervised) 18.81 26.81 10.00
LLM2Vec(Supervised) 41.45 40.33 21.05
SFR-Embedding-2_R 42.18 41.34 24.69
gte-Qwen2-7B-instruct 45.98 40.60 23.48
LLaMA2-inbatch-M 41.67 34.19 16.15
AutoRegEmbed 42.49 38.16 19.79
Table 2: Evaluation results on MS MARCO, NFcorpus, and SCIDOCS.
@@ 4.4 Ablation Study
To verify the effectiveness of AutoRegEmbed, we conducted an ablation study. First, we removed Conditional Distribution Alignment to evaluate its impact on model performance. Second, since Equation LABEL:eq7 was derived empirically in our previous work, we tested different variants of this equation to confirm that it remains the optimal choice. Different variants of Equation LABEL:eq7 include Log_sigmoid, which maps similarity to a logarithmic scale for integration with the exponential function e, as well as KL divergence and JS divergence, which quantify the distance between the conditional probability distributions of positive and negative sample embeddings in distinct ways.
@@
Figure 3: We evaluate the learning efficiency of our method against traditional contrastive learning on 10 STS datasets, comparing their performance under the same number of samples. Further details are provided in Appendix A.
Table 3 presents the ablation results. The experiments on different tasks indicate that Conditional Distribution Alignment improves performance by 9.17%, while Information Compression contributes a 16.99% improvement, demonstrating the effectiveness of both tasks. Additionally, experiments on variants of Equation LABEL:eq7 reveal that, although using a logarithmic scale for similarity and employing KL or JS divergence to measure distribution distance are more intuitive approaches, they do not surpass the performance of the original loss function in Equation LABEL:eq7. Thus, Equation LABEL:eq7 can be regarded as a more effective loss function. The specific equations and more detailed analysis are provided in Appendix B.
@@ 4.5 Learning Efficiency
To verify that AutoRegEmbed is better suited for LLMs, we compare its performance with four contrastive learning baselines under the same training data. Figure 3 shows that as the training data increases, the performance of both AutoRegEmbed and other contrastive learning methods improves, but AutoRegEmbed exhibits the fastest growth. Notably, with just 15,000 samples, AutoRegEmbed already surpasses the maximum performance of other contrastive learning models. The results indicate that AutoRegEmbed significantly outperforms the baseline models in learning efficiency.
@@ 6 Limitations
To mitigate this risk, we recommend that users of AutoRegEmbed ensure that the training data is carefully curated, and ideally, filtered for harmful content. Additionally, users should be cautious when applying AutoRegEmbed to sensitive domains, where the generation of unsafe or biased embeddings could have significant consequences.
Acknowledgments
This work was supported by the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the National Natural Science Foundation of China (NSFC) under Grants No. 62276248, the Key Research and Development Program of Xinjiang Uyghur Autonomous Region Grant No. 2024B03026, the Beijing Nova Program under Grants No. 20250484765, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.
@@ Appendix A Implementation details
For the information compression task, we set the learning rate to 2e-5, the batch size to 32, and train for 2 epoch. To represent the semantics of the input, we use 5 compressed tokens. For the conditional distribution alignment task, the learning rate is set to 5e-6, with a batch size of 32 and 4 epochs. The temperature parameters $\tau$ and $\beta$ are set to 0.05 and 0.1. For the above two tasks, we set the maximum token length of context, instruction, and target to 512. Furthermore, we employ the bfloat16 format, enable FlashAttention 2, and train on four A100-80G GPUs with DeepSpeed and Zero-2. The information compression task takes 20 minutes, while the conditional distribution alignment task, involving 50,000 samples, takes approximately 1 hour. In theory, our approach eliminates the need for in-batch negative sampling, which substantially reduces memory usage. In traditional contrastive learning, a batch size of 64 requires 64 * 64=4096 pairwise computations. In contrast, our method requires only 64, reducing the number of pairwise comparisons per step by a factor of 64 and lowering memory consumption accordingly.
@@ Appendix B Variants of Equation LABEL:eq7
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})}).$
@@
$S_{2}(q,d^{-}_{i})=-\sigma(\frac{1}{T}\sum_{t=1}^{T}\mathrm{KL}(p_{\theta_{E}}(d^{+}_{t}|d^{-}_{i,<t},e_{d^{-}_{i},I_{\mathrm{self}}}),$
@@
$S_{2}(q,d^{-}_{i})=-\sigma(\frac{1}{T}\sum_{t=1}^{T}\mathrm{JS}(p_{\theta_{E}}(d^{+}_{t}|d^{-}_{i,<t},e_{d^{-}_{i},I_{\mathrm{self}}}),$
@@
Why do KL divergence and JS divergence fail to achieve good performance in this setting?
We hypothesize that the effectiveness of using log-odds ratios stems from its ability to directly supervise the generation probability of a specific label token, thus providing a strong and targeted learning signal. In contrast, KL and JS divergence operate over the entire output distribution across the vocabulary. While these measures are theoretically well-founded for capturing distributional differences, they often suffer from instability in practice—especially in models like LLaMA2, which have large vocabularies ( 30,000 tokens). This instability weakens the gradient signal and ultimately degrades performance.
Method BIOSSES SICK-R STS12 STS13 STS14 STS15 STS16 STS17 STS22 STS-B Avg.
AutoRegEmbed-LLaMA2(Tanh) 83.97 80.75 80.58 86.92 83.19 88.98 86.96 85.80 65.91 85.98 84.77(7) / 82.90(10)
AutoRegEmbed-LLaMA2(Sigmoid) 85.50 79.07 79.57 86.90 83.28 88.45 86.57 88.61 66.16 86.59 84.35(7) / 83.07(10)
Table 4: Performance comparison of different $\sigma$ functions for AutoRegEmbed under the same settings.
Retrieval Task
$I_{\mathrm{next}}$
Use one word to represent the query in a
retrieval task. The word is: “
$I_{\mathrm{self}}$
Use one word to represent the passage in a
retrieval task. The word is: “
STS Task
$I_{\mathrm{next}}$ This sentence means in one word: “
$I_{\mathrm{self}}$ This sentence means in one word: “
Table 5: Instructions used for Retrieval and STS tasks.
Appendix C Selection Analysis of the $\sigma$
To stabilize exponential operations, we apply the $\sigma$ function to map distribution-based distances to a bounded range. Sigmoid and Tanh are typical choices. Table 4 presents their comparative performance under uniform settings.
Overall, under identical data and training settings, the Tanh variant slightly outperforms Sigmoid on 7 STS datasets, while Sigmoid shows a marginal advantage across all 10 datasets. Given the minimal performance gap, either function is a reasonable choice.
Appendix D Explanation of $I_{\mathrm{next}}$ and $I_{\mathrm{self}}$
Embedding tasks can be broadly categorized into sentence-to-sentence (STS) and sentence-to-document (retrieval), each associated with distinct instructions, $I_{\mathrm{next}}$ and $I_{\mathrm{self}}$. Based on prior studies (Jiang et al., 2024), the instructions used in this work are summarized in Table 5. Due to the symmetric nature of STS, both instructions are the same.
Appendix E Analysis of Alignment Strategies for Conditional Probability Distributions
The motivation behind our naive alignment strategy is as follows: $I_{\mathrm{next}}$ aims for $q$ to generate $d$, while $I_{\mathrm{self}}$ guides $d$ to generate itself. Therefore, we align the probability distribution of $q$ generating $d^{+}$ with that of $d^{+}$ generating itself. Intuitively, when the conditional distributions produced by their embeddings $e_{q,I_{\mathrm{next}}}$ and $e_{d^{+},I_{\mathrm{self}}}$ are closely matched, their embeddings are likely to be similar as well. Additionally, inspired by Direct Preference Optimization (DPO), we introduce negative sample feedback by encouraging $e_{q,I_{\mathrm{next}}}$ to increase the distance to negative samples while decreasing the distance to positive ones.
However, the use of instructions enables more sophisticated strategies for aligning the probability distributions between $q$ and $d$. To explore this, we conducted a comprehensive and detailed analysis. We begin by introducing several alignment strategy configurations.
Strategy 1
In this strategy, we align the conditional probability distribution of generating $q$ rather than $d^{+}$. Specifically, we adjust the instructions to treat $q$ as the anchor, aligning the probability distribution between $q$ and $d^{+}$. This is implemented by simply replacing the corresponding prompts, as detailed in Equation LABEL:eq11. For the STS task, where symmetry holds, $I_{\mathrm{self}}$ and $I_{\mathrm{prev}}$ are identical. For the asymmetric retrieval task, an example of $I_{\mathrm{prev}}$ could be: “Use a word to express the question that this article can answer:“”.
$S_{1}(q,d^{+})=-\sigma(\beta\ |\mathrm{log}\frac{p_{\theta_{E}}(q|e_{q,I_{\mathrm{self}}})}{p_{\theta_{E}}(q|e_{d^{+},I_{\mathrm{prev}}})}|)$ (3)
$S_{2}(d^{+},d^{-}_{i};q)=-\sigma(\beta\ \mathrm{log}\frac{p_{\theta_{E}}(q|e_{d^{+},I_{\mathrm{prev}}})}{p_{\mathrm{ref}}(q|e_{d^{+},I_{\mathrm{prev}}})}$
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(q|e_{d^{-}_{i},I_{\mathrm{prev}}})}{p_{\mathrm{ref}}(q|e_{d^{-}_{i},I_{\mathrm{prev}}})}).$
Strategy 2
We treat both $q$ and $d^{+}$ as anchors and compute a weighted similarity between their respective conditional distributions.
$S_{1}(q,d^{+})=-\frac{1}{2}\sigma(\beta\ |\mathrm{log}\frac{p_{\theta_{E}}(q|e_{q,I_{\mathrm{self}}})}{p_{\theta_{E}}(q|e_{d^{+},I_{\mathrm{prev}}})}|)$ (4)
$-\frac{1}{2}\sigma(\beta\ |\mathrm{log}\frac{p_{\theta_{E}}(d^{+}|e_{q,I_{\mathrm{next}}})}{p_{\theta_{E}}(d^{+}|e_{d^{+},I_{\mathrm{self}}})}|)$
$S_{2}(d^{+},d^{-}_{i};q)=-\frac{1}{2}\sigma(\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{+}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{+}|e_{q,I_{\mathrm{next}}})}$
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})})$
$-\frac{1}{2}\sigma(\beta\ \mathrm{log}\frac{p_{\theta_{E}}(q|e_{d^{+},I_{\mathrm{prev}}})}{p_{\mathrm{ref}}(q|e_{d^{+},I_{\mathrm{prev}}})}$
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(q|e_{d^{-}_{i},I_{\mathrm{prev}}})}{p_{\mathrm{ref}}(q|e_{d^{-}_{i},I_{\mathrm{prev}}})}).$
Method Params BIOSSES SICK-R STS12 STS13 STS14 STS15 STS16 STS17 STS22 STS-B Avg.
Strategy 1 7B 86.31 81.97 81.05 86.27 82.80 88.51 86.31 86.43 64.23 86.35 84.75(7) / 83.02(10)
Strategy 2 7B 83.34 81.87 79.82 84.11 81.97 89.11 86.39 88.79 66.11 86.90 84.32(7) / 82.84(10)
Strategy 3 7B 86.03 83.20 78.35 81.54 81.27 87.80 87.04 89.16 64.55 86.28 83.64(7) / 82.52(10)
Strategy 4 7B 84.10 80.74 78.09 83.79 81.77 88.01 86.14 89.64 64.14 87.16 83.67(7) / 82.36(10)
AutoRegEmbed 7B 85.50 79.07 79.57 86.90 83.28 88.45 86.57 88.61 66.16 86.59 84.35(7) / 83.07(10)
Table 6: Performance comparison of AutoRegEmbed under different alignment strategies on STS benchmarks.
Strategy 3
We can flexibly expand the range of negative samples by varying instructions—for example, by increasing the probability of positive samples generating themselves, while decreasing the probability of them generating negative samples.
$S_{2}(d^{+},d^{-}_{i};q)=-\frac{1}{2}\sigma(\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{+}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{+}|e_{q,I_{\mathrm{next}}})}$
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})})$
$-\frac{1}{2}\sigma(\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{+}|e_{d^{+},I_{\mathrm{self}}})}{p_{\mathrm{ref}}(d^{+}|e_{d^{+},I_{\mathrm{self}}})}$
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}_{i}|e_{d^{+},I_{\mathrm{self}}})}{p_{\mathrm{ref}}(d^{-}_{i}|e_{d^{+},I_{\mathrm{self}}})})$
Strategy 4
Similar to Strategy 3, we can also increase the probability of negative samples generating themselves, while reducing the probability of them generating positive samples, as shown in Equation LABEL:eq14.
Method Params BIOSSES SICK-R STS12 STS13 STS14 STS15 STS16 STS17 STS22 STS-B Avg.
AutoRegEmbed($\tau{=}0.1,\beta{=}0.1$) 7B 85.50 79.07 79.57 86.90 83.28 88.45 86.57 88.61 66.16 86.59 84.35(7) / 83.07(10)
AutoRegEmbed($\tau{=}0.02,\beta{=}0.1$) 7B 84.90 79.85 78.58 85.54 84.64 88.71 86.88 87.57 65.61 86.02 84.32(7) / 82.83(10)
AutoRegEmbed($\tau{=}0.05,\beta{=}0.1$) 7B 84.65 81.46 79.98 86.35 83.33 89.21 86.91 87.67 65.90 86.98 84.89(7) / 83.24(10)
AutoRegEmbed($\tau{=}0.2,\beta{=}0.1$) 7B 83.85 81.72 79.77 86.73 83.19 88.41 86.53 87.69 66.22 86.37 84.67(7) / 83.05(10)
AutoRegEmbed($\tau{=}1.0,\beta{=}0.1$) 7B 81.23 80.57 77.63 83.90 81.92 87.08 85.75 88.18 63.91 85.95 83.26(7) / 81.61(10)
AutoRegEmbed($\tau{=}0.1,\beta{=}0.2$) 7B 84.06 81.49 80.32 87.15 83.49 88.67 86.85 87.22 66.45 86.56 84.93(7) / 83.23(10)
AutoRegEmbed($\tau{=}0.1,\beta{=}0.3$) 7B 84.50 79.56 79.45 86.62 83.55 87.78 86.01 89.77 65.63 86.13 84.16(7) / 82.90(10)
AutoRegEmbed($\tau{=}0.1,\beta{=}0.4$) 7B 84.27 81.04 78.92 85.76 82.44 88.54 86.05 87.48 65.99 86.08 84.12(7) / 82.65(10)
AutoRegEmbed($\tau{=}0.05,\beta{=}0.2$) 7B 84.31 79.65 79.59 84.16 81.95 89.53 87.54 89.37 66.78 87.48 84.27(7) / 83.04(10)
Table 7: Ablation of temperature $\tau$ and alignment weight $\beta$ on STS benchmarks. The Avg. column shows results over 7 and 10 datasets.
Table 6 presents the performance of the LLaMA2-7B model trained on the STS split of the MEDI dataset under various alignment strategies. Both $\tau$ and $\beta$ are fixed at 0.1.
$S_{2}(d^{+},d^{-}_{i};q)=-\frac{1}{2}\sigma(\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{+}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{+}|e_{q,I_{\mathrm{next}}})}$
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})})$
$-\frac{1}{2}\sigma(\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}_{i}|e_{d^{-}_{i},I_{\mathrm{self}}})}{p_{\mathrm{ref}}(d^{-}|e_{d^{-}_{i},I_{\mathrm{self}}})}$
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{+}|e_{d^{-}_{i},I_{\mathrm{self}}})}{p_{\mathrm{ref}}(d^{+}|e_{d^{-}_{i},I_{\mathrm{self}}})})$
The experimental results reveal a consistent trend: the simpler the modification to the original loss function (Equation LABEL:eq7), the better the performance. Specifically, Strategy 1 alters the generation target from $d^{+}$ to $q$. Given the inherent symmetry between the two, this change leads to no significant performance difference. Strategy 2 computes a weighted average of the generation probabilities for $q$ and $d^{+}$; however, despite its increased computational cost, it does not improve performance. Strategies 3 and 4 introduce more complex alignment designs involving positive and negative samples, but these result in inferior performance.
Appendix F Analysis of Temperature Coefficients $\tau$ and $\beta$
The reasons for introducing these two temperature coefficients are as follows:
  • We introduce the temperature coefficient $\tau$ to align our loss function more closely with the InfoNCE objective. Intuitively, $\tau$ controls the strength of separation between positive and negative samples—the lower the value, the sharper the contrast enforced by the loss.
  • The temperature coefficient $\beta$ is inspired by the DPO (Direct Preference Optimization) loss, which uses a similar scaling factor to modulate the influence of preference differences—also expressed as log probability ratios. In our setting, $\beta$ controls the sensitivity of the loss to differences between distributions.
To empirically assess the impact of these hyperparameters, we conducted a grid search using the LLaMA2-7B model and the STS split from the MEDI dataset. For $\tau$, we adopted values commonly used in contrastive learning: 0.02, 0.05, 0.1, 0.2, 1.0{0.02,\ 0.05,\ 0.1,\ 0.2,\ 1.0}. For $\beta$, we selected values based on prior DPO studies: 0.1, 0.2, 0.3, 0.4{0.1,\ 0.2,\ 0.3,\ 0.4}. After systematic evaluation, we identified the best-performing $(\tau,\beta)$ pair.
To empirically assess the impact of these hyperparameters, we conducted a grid search using the LLaMA2-7B model and the STS split from the MEDI dataset. For $\tau$, we adopted values commonly used in contrastive learning: 0.02, 0.05, 0.1, 0.2, 1.0{0.02,\ 0.05,\ 0.1,\ 0.2,\ 1.0}. For $\beta$, we selected values based on prior DPO studies: 0.1, 0.2, 0.3, 0.4{0.1,\ 0.2,\ 0.3,\ 0.4}. After systematic evaluation, we observe that the model performs better when $\tau=0.05$ or $\beta=0.2$. Furthermore, except for a few extreme cases (e.g., $\tau=1.0$ or $\beta=0.4$), the model’s performance remains relatively stable across different settings, which demonstrates the robustness of our method. When $\tau=0.05$ and $\beta=0.2$ are applied simultaneously, the model does not exhibit improved performance. This suggests that the effects of these two temperature parameters may not be orthogonal.